OpenStudy (anonymous):

Write the equation of a parabola with a vertex at (0, 0) and a directrix y = -1.

OpenStudy (anonymous):

y = x^2 ?

OpenStudy (anonymous):

okay so, either y = -1/4x^2 or y = 1/4x^2 its -1/4x^2, eh?

OpenStudy (anonymous):

no, pretty sure it will be positive. since the directrix is -1

OpenStudy (anonymous):

whaat, really ? strange. i thought for sure it would be negative.

OpenStudy (anonymous):

honestly i don't have that much experience with directrix... but from what i googled, the directrix is equidistant from the vertex as the focus is, and the focus is inside the cup of the parabola.

OpenStudy (anonymous):

the directrix can't cross the parabola, and y = x^2 stays above the x axis

OpenStudy (anonymous):

so if you had y = -1/4x^2, the directrix would intersect with the parabola, which isn't allowed

OpenStudy (anonymous):

ohh alright. that makes sense then. so, would you be able to help me with another one super quick? cause i have this question ... Write the equation of a parabola with a vertex at (0, 0) and a focus at (-4, 0). and im unfamilar with how to do these as well. =/ once i get like an example i can usually work off of that. (:

OpenStudy (anonymous):

this one would be of the form y = -Ax^2 i don't know what A is, i'm sure there's some formula for finding it based off what the focus is

OpenStudy (anonymous):

huh. that doesnt help my case. but thanks! haha. ill just have to find an example quick i seeepose.

OpenStudy (anonymous):

is your focus (-4, 0) or (0, -4)?

OpenStudy (anonymous):

if the focus is (-4, 0) then your equation will be x = -(1/16)y^2 if the focus is (0, 4) then your equation will be y = -(1/16)x^2